OpenStudy (jordan123321):

What is the circumference of a circle with a radius of 0.8 m? A. m B. m C. m D. m

OpenStudy (anonymous):

What are the answers?

OpenStudy (haseeb96):

For circumference of circle we use this formula 2*pi*r here r=radius pi=3.14 now r= 0.8m (given in the question) now just put this value of r and pi in the above formula and get the answer

OpenStudy (haseeb96):

and tell me what u got?

OpenStudy (anonymous):

Which equation is represented by the table? x y (x, y) 0 3 (0, 3) 1 5 (1, 5) 2 7 (2, 7) A. y = 3x + 2 B. y = 5x C. y = x + 3 D. y = 2x + 3

OpenStudy (anonymous):

haseeb can i also try to answer this?

OpenStudy (anonymous):

C=2πr=2·π·0.8≈5.02655 C≈5.03